Lesson Activity: A. A population consists of the numbers 2, 4, 8, 10 and 5. Let us list all the possible samples of size 3 from this population and construct the sampling distribution of the sample mean. B. Complete the statement by filling in the blank. Write your answer on a separate sheet of paper. A ___________ is a frequency distribution using the means computed from all possible random samples of a specific size taken from a population. To get the possible samples use the formula ______, where N is the ________ and n is the ____ size to be taken. The total probability of the sample mean must be equal to ____.
1
Expert's answer
2021-10-25T03:10:53-0400

Since the size of the population is "N=5" sample sizes of "n=3" are to be drawn from this population then, there are "\binom{N}{n}=\binom{5}{3}=5!/(3!*2!)=10" possible samples that can be drawn.

The possible samples and their respective sample means are given below.

The sample means are derived using the formula below,

"\bar{x}=\sum(x)/n" where "x" are the data values for each sample and "n=3" for each sample.


"sample" "\bar{x}"

2,4,8 (2+4+8)/3=4.67

2,4,10 (2+4+10)/3=5.33

2,4,5 (2+4+5)/3=3.67

2,8,10 (2+8+10)/3=6.67

2,8,5 (2+8+5)/3=5

2,10,5 (2+10+5)/3=5.67

4,8,10 (4+8+10)/3=7.33

4,8,5 (4+8+5)/3=5.67

4,10,5 (4+10+5)/3=6.33

8,10,5 (8+10+5)/3=7.67

Since we are drawing at random, each sample will have the same probability of being chosen. Therefore probabilities for these sample means are,

"p(\bar{x}=4.67)=1/10"

"p(\bar{x}=5.33)=1/10"

"p(\bar{x}=3.67)=1/10"

"p(\bar{x}=6.67)=1/10"

"p(\bar{x}=5)=1/10"

"p(\bar{x}=5.67)=1/5"

"p(\bar{x}=7.33)=1/10"

"p(\bar{x}=6.33)=1/10"

"p(\bar{x}=7.67)=1/10"


A "sampling\space distribution" is a frequency distribution using the means computed from all possible random samples of a specific size taken from a population. To get the possible samples use the formula "\binom{N}{n}=N!/(n!*(N-n)!)", where "N" is the "population\space size" and "n" is the "sample" size to be taken. The total probability of the sample mean must be equal to "1"